OpenStudy (anonymous):

What is the area of a circle with a diameter of 6?

OpenStudy (anonymous):

Area = pie * r*r

OpenStudy (anonymous):

R = D/2 = 6/2 = 3

OpenStudy (anonymous):

\[3.14*3*3 = 28.26\]

OpenStudy (anonymous):

these are the only answers i have to choose from 3π 6π 9π 36π

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

3*3 = 9 so 9 pie